OpenStudy (anonymous):

Find S15 for 1 + 1.5 + 2.25 + 3.375 +...

OpenStudy (anonymous):

seems more like r = ?

OpenStudy (anonymous):

r=1.5?

OpenStudy (solomonzelman):

another parthkohli ;) ?

OpenStudy (xguardians):

The Pattern is multiplying each one by 1.5.

OpenStudy (paki):

yeah... parthkohli (2) :D

OpenStudy (anonymous):

This is a geometric series with r = 1.5, as you mentioned. Try using the summation formula to find the \(15\)th term.

OpenStudy (anonymous):

*sigh* I made this account for some reasons. Anyway, yes, this is the gist of it.

OpenStudy (anonymous):

291.92?

OpenStudy (anonymous):

That's not exactly what I'm getting.

OpenStudy (anonymous):

ill try again

OpenStudy (anonymous):

so 873? :) :)

OpenStudy (anonymous):

I guess so.

OpenStudy (anonymous):

would you help me with another one pleasee??

OpenStudy (anonymous):

Surely.

OpenStudy (anonymous):

I need to find the three geometric means between 1/2 and 8

OpenStudy (anonymous):

There are many ways to do this - all of them similar. One way is to define a geometric progression with the first term as 1/2 and the fifth term as 8.

OpenStudy (anonymous):

ok :)?

OpenStudy (anonymous):

All geometric progressions are in the form,\[a, ~ar, ~ar^2 \cdots ar^{n -1}\]We know that are geometric progression has the first term 1/2 and the last term 8. The three "geometric means" are actually the middle three terms that are in the geometric progression. Our geometric progression is in the form\[1/2, ~~ r\cdot 1/2 , ~~ r^2 \cdot 1/2 , ~ ~ r^3 \cdot 1/2, ~ ~ r^4 \cdot 1/2\]

OpenStudy (anonymous):

But we know for a fact that the fifth term is \(8\). The fifth term is in the form \(r^4 \cdot 1/2\), which means\[r^4 \cdot \dfrac{1}{2} = 8 \implies r^4 = 16 \implies r = 2\]

OpenStudy (anonymous):

Now that you know \(r\), you can easily calculate the three terms. Do you follow?

OpenStudy (anonymous):

yes :)?

OpenStudy (anonymous):

OK - what do you get?

OpenStudy (anonymous):

1, 2, 4?

OpenStudy (anonymous):

That's right!

OpenStudy (anonymous):

THANKYOUUU!!!!

OpenStudy (anonymous):

No problem! And my main account is @ParthKohli - not this one, lol.

OpenStudy (anonymous):

already a fan :) thanks!